OpenStudy (anonymous):

What is the simplified expression for –3cd – d(2c – 4) – 4d? –5cd cd – 8d 5cd – 8d –5cd – 8d

OpenStudy (cwrw238):

be careful with the signs when you distribute -d over (2c - 4)

OpenStudy (anonymous):

thank you but I still need help

OpenStudy (cwrw238):

-d(2c - 4) = -d*2c -d*-4 = -2cd + 4d (note -d * -4 = + 4d)

OpenStudy (cwrw238):

so you get -3cd - 2cd +4d - 4d can you continue?

OpenStudy (anonymous):

yeah I got that

OpenStudy (cwrw238):

great

OpenStudy (anonymous):

now I am lost

OpenStudy (cwrw238):

well its important to rememger that you can only add/subtract LIKE terms eg x and 3x are like terms so are 2y and - 3y here we have -3cd and -2cd and -4d and + 4d what does -3cd - 2cd simplify to?

OpenStudy (cwrw238):

can you work out - 3 - 2 ?

OpenStudy (anonymous):

-5

OpenStudy (cwrw238):

right so -3cd - 2cd = -5cd

OpenStudy (cwrw238):

and what does -4 + 4 come to?

OpenStudy (anonymous):

0

OpenStudy (cwrw238):